# Some new findings concerning value distribution of a pair of delay-differential polynomials

### v1

@@ 1. Introduction
From (2) and (2), we see the small functions of $f_{1}$ and $f_{2}$ are also the the small functions of $f$ and $g$, therefore $p_{1}$ and $p_{2}$ are small functions of $f$ and $g$. Using the same methods of Case 1 to (3.6) and (3.7), we also can get $N(r,\frac{1}{f})=S(r,f)$ and $N(r,\frac{1}{g})=S(r,g)$. By the Weierstrass’s factorization theorem, we can get $f(z)=A_{1}(z)e^{B_{1}(z)}$, where $A_{1}$ and $B_{1}$ are entire function, such that $T(r,A_{1})=S(r,f)$, $g(z)=A_{2}(z)e^{B_{2}(z)}$, where $A_{2}$ and $B_{2}$ are entire function, such that $T(r,A_{2})=S(r,f)$. Let $\frac{L(z,f)}{f(z)}=a_{1}$, By applying the logarithmic derivative lemma, we can obtain $m(r,\frac{L(z,f)}{f})\leq S(r,f)$. Since $N(r,\frac{1}{f})=S(r,f)$, we have $T(r,\frac{L(z,f)}{f})=m(r,\frac{L(z,f)}{f})+N(r,\frac{L(z,f)}{f})=S(r,f)$. Thus $a_{1}$ is a small function of $f$. Using the same derivation, we get $L(z,g)=g(z)a_{2}$, where $a_{2}$ is a small function of $g$. Substituting the expressions for $f$, $L(z,f)$, $g$ and $L(z,g)$into the equation (3.6) and (3.7), we obtain
(3.8) $\begin{cases}a_{3}e^{nB_{1}+B_{2}}-a=p_{1}e^{b_{1}},\\ a_{4}e^{mB_{1}+B_{2}}-a=p_{2}e^{b_{2}},\end{cases}$
where $a_{3}$ and $a_{4}$ are small function of $f$ and $g$. Since $m(r,g)\leq m(r,f)$, we have $T(r,f_{1})=T(r,a_{3}e^{nB_{1}+B_{2}})\geq m(r,f)$. By the second fundamental theorem concerning three small functions, we have
$T(r,a_{3}e^{nB_{1}+B_{2}})<\overline{N}(r,\frac{1}{a_{3}e^{nB_{1}+B_{2}}})+\overline{N}(r,\frac{1}{a_{3}e^{nB_{1}+B_{2}}-a})+\overline{N}(r,{a_{3}e^{nB_{1}+B_{2}}}))+S(r,f)=S(r,f),$
which is impossible. Therefore at least one of $f^{n}L(z,g)-a$ and $g^{n}L(z,f)-a$ has infinite zeros. This means that $a(z)$ cannot be a common non-zero generalized Picard exceptional for both $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$.
@@
$m(r,g)\leq m(r,f)$ and $a$ is not a small function of $f_{2}$. We claim $f^{n}L(z,g)-a$ has infinite zeros. Otherwise, $f^{n}L(z,g)-a$ has finite zeros. As the same as the Case 1, we get equation (3.1)-(2) and $f(z)=A_{3}(z)e^{B_{3}(z)}$, where $A_{3}$ and $B_{3}$ are entire function, such that $T(r,A_{3})=S(r,f)$. Since $a$ is not a small function of $f_{2}$, then $T(r,f_{2})\leq O(T(r,a))=S(r,f)$.
By applying the logarithmic derivative lemma, we can obtain $m(r,\frac{L(z,f)}{f})\leq S(r,f)$. Since $N(r,\frac{1}{f})=S(r,f)$, we have $T(r,\frac{L(z,f)}{f})=m(r,\frac{L(z,f)}{f})+N(r,\frac{L(z,f)}{f})=S(r,f)$. Thus $\frac{L(z,f)}{f}=a_{5}$, where $a_{5}$ is small function of $f$. There fore $g^{m}=\frac{f_{1}}{L(z,f)}=f_{1}\frac{1}{a_{5}f}$, which means $g=A_{4}e^{\frac{-B_{3}}{m}}$, which $A_{4}$ is a small function of $f$. As the same as $L(z,f)$, we have $L(z,g)=a_{6}g$, where $a_{6}$ is a small function of $g$.
@@
$a_{7}e^{\frac{n-1}{m}B_{3}}-a=pe^{b}$
where $a_{7}$ is a small function of $f$. Since $n\geq 2$, we have $T(r,a_{7}e^{\frac{n-1}{m}B_{3}})=T(r,f)+S(r,f)$. By the second fundamental theorem concerning three small functions, we have
$T(r,a_{7}e^{\frac{n-1}{m}B_{3}})<\overline{N}(r,\frac{1}{a_{7}e^{\frac{n-1}{m}B_{3}}})+\overline{N}(r,\frac{1}{a_{7}e^{\frac{n-1}{m}B_{3}}-a})+\overline{N}(r,{a_{7}e^{\frac{n-1}{m}B_{3}}}))+S(r,f)=S(r,f),$
which is impossible. Therefore $f^{n}L(z,g)-a$ has infinite zeros. This means that $a(z)$ cannot be a common non-zero generalized Picard exceptional for both $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$.
@@
We will prove the Theorem 2.2 by using the idea of prove Theorem 2.1 in two cases.
@@
Suppose $\min\{m,n\}\geq 2$ and $f^{n}L(z,g)$ and $g^{m}L(z,f)$ have simultaneously non-zero Picard exceptional small functions. By Weierstrass’s factorization theorem, we assume that
@@
If $g$ is a small function of $f$. Since $a_{1}$, $p_{1}$ are small functions of $f^{n}L(z,g)$, then $a_{1}$ and $p_{1}$ are also a small functions of $f$. If $g$ is not a small function of $f$. Since $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f^{n}L(z,g)$ and $g^{m}L(z,f)$, receptively, it is easy to get $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f$ and $g$, receptively. These two situations are similar to case 1 and case 2.2 of Theorem 2.1, then we can also derive a contradiction through the same method. Therefore we omit the prove.
@@
Suppose $f^{n}g$ and $gf$ have simultaneously non-zero Picard exceptional small functions. By Weierstrass’s factorization theorem, we assume that
@@
If $g$ is a small function of $f$, we can derive a contradiction through the same method of Case 1. Therefore, we only need to consider $g$ is not a small function of $f$. Thus $T(r,f)=O(T(r,g))$, $S(r,f)=S(r,g)$. Since $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f^{n}L(z,g)$ and $g^{m}L(z,f)$, receptively, it is easy to get $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f$ and $g$, receptively.
Since $a_{1}$ is a small function of $f^{n}L(z,g)$, then $a_{1}$ is also a small function of $f$. Using the same methods of Case 1 of the prove of Theorem 2.1 to the first equation (3.10) , we also an get $N(r,\frac{1}{f})=S(r,f)$. By the Weierstrass’s factorization theorem, we can get $f(z)=A(z)e^{B(z)}$, where $A$ is a small function of $f$, $B$ is an entire function. Substituting the expression for $f$ into the second equation of (3.10), we get
(3.11) $g=A_{2}e^{b_{2}(z)-B(z)}+A_{3}e^{-B(z)},$
where $A_{2}$, $A_{3}$ are small function of $f$ and $g$. Substituting the expressions for $f$ and $g$ into the first equation of (3.10), we get

### v3

@@ 2. Hayman’s conjecture on a pair of delay-differential polynomials
From (2) and (2), we see the small functions of $f_{1}$ and $f_{2}$ are also the the small functions of $f$ and $g$, therefore $p_{1}$ and $p_{2}$ are small functions of $f$ and $g$. Using the same methods of Case 1 to (3.6) and (3.7), we also can get $N(r,\frac{1}{f})=S(r,f)$ and $N(r,\frac{1}{g})=S(r,g)$. By the Weierstrass’s factorization theorem, we can get $f(z)=A_{1}(z)e^{B_{1}(z)}$, where $A_{1}$ and $B_{1}$ are entire function, such that $T(r,A_{1})=S(r,f)$, $g(z)=A_{2}(z)e^{B_{2}(z)}$, where $A_{2}$ and $B_{2}$ are entire function, such that $T(r,A_{2})=S(r,f)$. Let $\frac{L(z,f)}{f(z)}=a_{1}$, By applying the logarithmic derivative lemma, we can obtain $m(r,\frac{L(z,f)}{f})\leq S(r,f)$. Since $N(r,\frac{1}{f})=S(r,f)$, we have $T(r,\frac{L(z,f)}{f})=m(r,\frac{L(z,f)}{f})+N(r,\frac{L(z,f)}{f})=S(r,f)$. Thus $a_{1}$ is a small function of $f$. Using the same derivation, we get $L(z,g)=g(z)a_{2}$, where $a_{2}$ is a small function of $g$. Substituting the expressions for $f$, $L(z,f)$, $g$ and $L(z,g)$ into the equations (3.6) and (3.7), we obtain
(3.8) $\begin{cases}a_{3}e^{nB_{1}+B_{2}}-\alpha=p_{1}e^{b_{1}},\\ a_{4}e^{mB_{1}+B_{2}}-\alpha=p_{2}e^{b_{2}},\end{cases}$
where $a_{3}$ and $a_{4}$ are small function of $f$ and $g$. By the second fundamental theorem concerning three small functions, we have
$T(r,a_{3}e^{nB_{1}+B_{2}})$ $<\overline{N}(r,\frac{1}{a_{3}e^{nB_{1}+B_{2}}})+\overline{N}(r,\frac{1}{a_{3}e^{nB_{1}+B_{2}}-\alpha})+\overline{N}(r,{a_{3}e^{nB_{1}+B_{2}}}))+S(r,f)$
$=S(r,f),$
which is impossible. Therefore at least one of $f^{n}L(z,g)-\alpha$ and $g^{n}L(z,f)-\alpha$ has infinitely many zeros. This means that $\alpha$ cannot be a common non-zero generalized Picard exceptional small function for both $f^{n}L(z,g)$ and $g^{m}L(z,f)$.
@@
$m(r,g)\leq m(r,f)$ and $\alpha$ is not a small function of $f_{2}$. We claim $f^{n}L(z,g)-\alpha$ has infinitely many zeros. Otherwise, $f^{n}L(z,g)-\alpha$ has finitely many zeros. As the same as the Case 1, we get equations (3.1)-(2) and $f(z)=A_{3}(z)e^{B_{3}(z)}$, where $A_{3}$ and $B_{3}$ are entire function, such that $T(r,A_{3})=S(r,f)$. Since $\alpha$ is not a small function of $f_{2}$, then $T(r,f_{2})\leq O(T(r,\alpha))=S(r,f)$.
By applying the logarithmic derivative lemma, we can obtain $m(r,\frac{L(z,f)}{f})\leq S(r,f)$. Since $N(r,\frac{1}{f})=S(r,f)$, we have $T(r,\frac{L(z,f)}{f})=m(r,\frac{L(z,f)}{f})+N(r,\frac{L(z,f)}{f})=S(r,f)$. Thus $\frac{L(z,f)}{f}=a_{5}$, where $a_{5}$ is a small function of $f$. Therefore $g^{m}=\frac{f_{1}}{L(z,f)}=f_{1}\frac{1}{a_{5}f}$, which means $g=A_{4}e^{\frac{-B_{3}}{m}}$, which $A_{4}$ is a small function of $f$. As the same as $L(z,f)$, we have $L(z,g)=a_{6}g$, where $a_{6}$ is a small function of $g$.
@@
$a_{7}e^{\frac{n-1}{m}B_{3}}-\alpha=pe^{b},$
where $a_{7}$ is a small function of $f$. Since $n\geq 2$, we have $T(r,a_{7}e^{\frac{n-1}{m}B_{3}})=O(T(r,f))+S(r,f)$. By the second fundamental theorem concerning three small functions, we have
$T(r,a_{7}e^{\frac{n-1}{m}B_{3}})<\overline{N}(r,\frac{1}{a_{7}e^{\frac{n-1}{m}B_{3}}})+\overline{N}(r,\frac{1}{a_{7}e^{\frac{n-1}{m}B_{3}}-\alpha})+\overline{N}(r,{a_{7}e^{\frac{n-1}{m}B_{3}}}))+S(r,f)=S(r,f),$
which is impossible. Therefore $f^{n}L(z,g)-\alpha$ has infinitely many zeros. This means that $\alpha$ cannot be a common non-zero generalized Picard exceptional small function for both $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$.
@@
We will prove the Theorem 2.2 by using similar ideas in proving Theorem 2.1, we divide the proof into two cases.
@@
Suppose $\min\{m,n\}\geq 2$ and $f^{n}L(z,g)$ and $g^{m}L(z,f)$ have simultaneously non-zero generalized Picard exceptional small functions. By Weierstrass’s factorization theorem, we assume that
@@
If $g$ is a small function of $f$. Since $a_{1}$, $p_{1}$ are small functions of $f^{n}L(z,g)$, then $a_{1}$ and $p_{1}$ are also small functions of $f$. If $g$ is not a small function of $f$. Since $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f^{n}L(z,g)$ and $g^{m}L(z,f)$, receptively, it is easy to get $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f$ and $g$, receptively. These two situations are similar to case 1 and case 2 of Theorem 2.1, then we can also derive a contradiction through the same method. Therefore we omit the proof.
@@
Suppose $f^{n}g$ and $gf$ have simultaneously non-zero generalized Picard exceptional small functions. By Weierstrass’s factorization theorem, we assume that
@@
If $g$ is a small function of $f$, we can derive a contradiction through the same method of Case 1. Therefore, we only need to consider $g$ is not a small function of $f$. Thus $T(r,f)=O(T(r,g))$, $S(r,f)=S(r,g)$. Since $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f^{n}L(z,g)$ and $g^{m}L(z,f)$, receptively, it is easy to get $a_{1}$, $p_{1}$ and $a_{2}$, $p_{2}$ are small functions of $f$ and $g$.
Since $a_{1}$ is a small function of $f^{n}L(z,g)$, then $a_{1}$ is also a small function of $f$. Using the same methods of Case 1 of the proof of Theorem 2.1 to the first equation (3.10), we also get $N(r,\frac{1}{f})=S(r,f)$. By the Weierstrass’s factorization theorem, we can get $f(z)=A(z)e^{B(z)}$, where $A$ is a small function of $f$, $B$ is an entire function. Substituting the expression for $f$ into the second equation of (3.10), we get
(3.11) $g(z)=A_{2}(z)e^{b_{2}(z)-B(z)}+A_{3}(z)e^{-B(z)},$
where $A_{2}$, $A_{3}$ are small functions of $f$ and $g$. Substituting the expressions for $f$ and $g$ into the first equation of (3.10), we get